OpenStudy (anonymous):

I cannot do this. It says consider a sequence generated by the formula f(n)=6(n)-4 starting with n=1. Generate the terms f(1),f(2),f(3)...

OpenStudy (anonymous):

f(1) = 6(1) - 4 = 2 f(2) = 6(2) - 4 = 8 f(3) = 6(3) - 4 = 14 So f(1), f(2), f(3) ... = 2, 8, 14 ...

OpenStudy (anonymous):

thou so so much we need help again if you can you are so nice

OpenStudy (anonymous):

My keyboard is broken. It meant to say thank you so so much

OpenStudy (anonymous):

consider the sequence given by the formula f(n)= 13n-1 starting with n=1, generate the first 5 terms of the sequence